OpenStudy (anonymous):

evaluate (c -b)2 + a2 for a= -2, b= -1, and c=7 i need help!!

OpenStudy (aum):

Are you sure it is (c -b)2 = a2 and not (c -b)2 + a2 ? (plus in the place of equal sign)

OpenStudy (anonymous):

yes. im sorry xD it was hard to write out the problem xD

OpenStudy (aum):

Just plug in the numbers: (c - b)^2 + a^2 for a= -2, b= -1, and c=7 (7 - -1)^2 + (-2)^2 = ?